# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ 4.1 The Non-Erasing Cases
In the non-erasing case, an interesting observation regarding $r_{ab}$ can be found. Reconstructing the reduction from [BREMER201215] in the setting of relational pattern languages, it can be shown that the inclusion problem for non-erasing relational pattern languages under $r_{ab}$ is actually also undecidable, even though $r_{ab}$ generally implies a significantly less restrictive behavior than $r_{=}$.
Proposition 5(✂)
Let $R:=\{r_{ab}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$. It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
Proof
This can be shown by reducing the problem of deciding whether the universal Turing machine $U$ has any valid computations to the problem of whether the NE-relational pattern language of one relational pattern under abelian equivalence is included in the NE-relational pattern language of another relational pattern under abelian equivalence. It is done by constructing two such relational patterns $(\alpha,\alpha_{R})$ and $(\beta,\beta_{R})$ for which we have that $L_{NE}(\alpha,R)\subseteq L_{NE}(\beta,\beta_{R})$ if and only if $U$ has no valid computation.
As mentioned above, this result follows from the same reduction as in [BREMER201215] for the inclusion problem of NE-pattern languages over the equality relation. Only some technical modifications have to be made to formally get it to work in the setting of relational pattern languages under $r_{ab}$. The proof covering the technical adaptions can be found in Appendix 0.B.4 for reference. ∎
As the construction used to show Proposition 5 only related variables to either enforce equality between length 1 substitutions (letters) or equality between unary strings, it also not only translates to the case of $r_{ab}$, but also to the reversal relation $r_{rev}$. Indeed, the reversal of a unary word $w$ is $w$ itself and the reversal of a letter $\mathtt{a}_{i}$ is clearly also just the letter $\mathtt{a}_{i}$. So, we can enforce the same behavior to select only one $x_{i,1}$ to become the letter #\# in valid substitutions while all other $x_{j,i}$ take the letter 0, as $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are always substituted only to unary words that are powers of the letter 0. With that, the $v$ parts will look exactly the same. Further, the counter values that are represented in simple predicates with the variables $\hat{x}_{i,j}$ will also only take the shape of unary words, as do the variables in the third predicate $\pi_{3}$ in the remaining cases not covered by the predicates $\pi_{1}$ and $\pi_{2}$. So, using an analogous construction as for Proposition 5, we obtain the following.
Corollary 2
Let $R:=\{r_{rev}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$. It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
@@ 4.2 The Erasing Cases
We continue with the additional cases for erasing relational pattern languages. First, it can be shown that the inclusion problem under the reversal relation $r_{rev}$ is also undecidable. We can use the same framework as in the result from Freydenberger and Reidenbach [Freydenberger2010] that shows undecidability in the erasing case under $r_{=}$, for all alphabets $\Sigma$ with $|\Sigma|\geq 2$, as a starting point, but we need to adapt it at several points in a more involved way to get the following theorem. Similarly to Proposition 5, in Appendix 0.B.5, we provide, for reference and completeness, a full reconstruction of the proof from [Freydenberger2010] in the framework of relational pattern languages with adaptations where they are necessary for the proof to work for $r_{rev}$.
Theorem 4.1(✂)
Let $R=\{r_{rev}\}$. For all alphabets $\Sigma$ with $|\Sigma|\geq 2$, given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is generally undecidable whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
Next, as for the non-erasing case, we can do something similar to show undecidability of the inclusion problem also relational pattern languages under abelian equivalence $r_{ab}$.
Theorem 4.2
Let $\Sigma$ be a finite alphabet with $|\Sigma|\geq 2$ and let $R=\{r_{ab}\}$. Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
Proof
This proof is analogous to the one of Theorem˜4.1. Therefore, we only describe how to get to the analogy and refer for a detailed proof to Theorem˜4.1. Let $A$ be a nondeterministic 2-counter automaton. Consider
$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$ and
with, for all $i\in\{1,\ldots,\mu\}$, and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{4}},x_{1_{5}},\ldots$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all $\eta_{i}\in X^{*}$ are terminal-free patterns for some $\mu$ and some $\hat{\beta}_{i}$ as in Theorem˜4.1. We have $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ with distinct variables $z_{i},z_{i^{\prime}},\hat{z}_{i_{1}},\hat{z}_{i_{2}},\hat{z}_{i_{3}},\hat{z}_{i_{4}},\hat{z}_{i_{5}}$ and $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
The main concept of the proof of Theorem˜4.1 is to show that iff $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{a},r_{\beta_{A}})$ then $A$ has no accepting computation. All possible non accepting computations are listed in the $\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}$ part where each $\hat{\beta}_{i},i\in\{1,\ldots,\mu\}$ encodes one non accepting computation together with selecting variables $x_{i_{j}}$ for $j\in{1,2,3}$. The part of $\beta_{A}$ and $\alpha_{A}$ before the first $^{6}$ ensures that exactly one pair $x_{i},x_{i}^{\prime},i\in\{1,\ldots,\mu\}$ is selected, resulting in that exactly one non accepting computation is selected and that exactly one $\ddot{\beta}_{i},i\in\{1,\ldots,\mu\}$ is selected. With this, we get that $\eta_{i},i\in\{1,\ldots,\mu\}$ matches $u$ and thus, that $z_{i}=0$ and $\hat{z}_{i_{1}}=\#$. We use these two variables to encode the 0 and the #\# in the non accepting computation encodings. If we can show that this selecting process also works in the setting $R=r_{ab}$ and that we get two variables with which we can encode the letters 0 and #\#, then this proof also works in this setting. For two variables $x,y$, we have if $x=0$ (or x=#x=\#) and $(x,y)\in r_{ab}$, then $y=0$ (or y=#y=\# respectively). Thus, we can model with the relation $r_{ab}$ identical variables if we know a one letter substitution of one variable.
[…]
$x_{i}$ $x_{i^{\prime}}$ works?
0 0 $00\neq 0\#$
0#0\# 0#0\# or #0\#0 $0#$, $0\#,\ 0\#^{2}0\neq 0\#^{3}$
$0\#^{3}0$ $0\#^{3}0$ $0\#^{3}00\#^{3}0=v\ v$
Table 1: Possibilities for the first match with vvv\ v
Thus, the only possibility is that for exactly one $i\in\{1,\ldots,\mu\}$ the variables $x_{i}$ and $x_{i^{\prime}}$ get substituted with $v$ and all other variables $x_{j},x_{j^{\prime}}$ for $j\neq i,j\in\{1,\ldots,\mu\}$ get substituted with $\epsilon$. This together with the matching from vuvv\ u\ v to $\ddot{\beta}_{1}\ldots\ddot{\beta}_{\mu}$ implies that $\eta_{i}$ needs to match $u$. For this matching between $\eta_{i}=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ and $u=0\#^{5}0$, we get the following possibilities. Remember that $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
$z_{i}$ $\hat{z}_{i_{1}}$ works?
$\epsilon$ 0 or #\# $0^{5},\ \#^{5}\neq 0\#^{5}0$
0 #\# $0\#^{5}0=0\#^{5}0=u$
0#0\# #\# $0\#\#^{5}\#0\neq 0\#^{5}0$
0##0\#\# #\# $0\#^{2}\#^{5}\#^{2}0\neq 0\#^{5}0$
0###0\#\#\# $\epsilon$ $0\#^{6}0\neq 0\#^{5}0$
Table 2: Possibilities for the match with $u$
Thus, the only possibility is that $z_{i}$ and $z_{i^{\prime}}$ get substituted with 0 and $\hat{z}_{i_{k}},k\in\{1,2,3,4,5\}$ get substituted with #\#. With this, we can encode the 0 and the #\# and thus, the proof of Theorem˜4.1 works in this setting analogously. ∎
This concludes the results regarding the inclusion problem for relational pattern languages covered in this paper.
4.3 Discussion and Challenges Regarding the Other Relations
Currently, we know that for relational patterns under the relations $r_{=}$, $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ the inclusion problem is undecidable. This mainly comes from the fact that either the relations can be used to simulate the behavior of $r_{=}$ generally (as in the case of $r_{ssq}$ and $r_{*}$) or it comes from the fact that essential properties used in relevant reductions can also be emulated using other relations (enforce equal letters or equal unary words using $r_{ab}$ or $r_{rev}$).
When it comes to the remaining relations $r_{|w|}$, $r_{perm}$, $r_{com^{*}}$, and $r_{com^{+}}$, we see that these properties are not trivially obtainable. The relations $r_{|w|}$ and $r_{perm}$ both do not force equal letters or equal unary words in alphabets that are at least of binary size. The relations $r_{com^{*}}$ and $r_{com^{+}}$ do not even force equal length between variables in valid substitutions, making, for example, the comparison of counter values impossible in the way it has been done in the previous reductions. Hence, it remains an worthwhile open question whether the inclusion problem becomes decidable at some point, for very structurally weak relations.
Question 1
Let $R=\{\tilde{r}\}$ for some relation $\tilde{r}\in\{r_{|w|},r_{perm},r_{com^{*}},r_{com^{+}}\}$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. Is the inclusion problem $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$, generally decidable?
5 The Equivalence Problem
In this section, we briefly discuss some results regarding the equivalence problem for non-erasing relational pattern languages. As we will see, we will obtain a series of small results that builds up to the main theorem of this section, stating that the equivalence problem for all non-erasing pattern languages over a binary equivalence relation that is antisymmetric on elements of length 1 (i.e., has equality for single letters) is decidable in polynomial time (covering the relations $\{r_{=},r_{ab},r_{con^{+}}\}$ from this paper’s discussion).
We define $r_{\alpha}]$ to be the transitive, reflexive and symmetric closure on some set of binary related variables $r_{\alpha}$. Additionally, for this section, we assume that variables are always introduced inside a pattern with increasing indices starting from $1$ (i.e., $x_{1}$, $x_{2}$, and so on). Due to space constraints, the technical proofs of almost all steps can be found in Appendix 0.C, while the steps themselves are supposed to give some intuition on why the main result of this section, Theorem 5.1, holds. We begin with a result regarding $r_{\alpha}]$ in the case of $r_{\alpha}$ being an equivalence relation.
Lemma 3
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation and let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. Then we have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$.
Proof
($\Rightarrow$): Let $w\in L_{NE}(\alpha,r_{\alpha})$. Then there exists $h\in H$ such that $h((\alpha,r_{\alpha}))=w$. Let $(\tilde{r},x,y)\in r_{\alpha}$ for $x,y\in\operatorname{var}(\alpha)$. Then $(h(x),h(y))\in\tilde{r}$ by definition. As $\tilde{r}$ is an equivalence relation, we also have $(h(y),h(x))\in\tilde{r}$, $(h(x),h(x))\in\tilde{r}$, and if there exists $(\tilde{r},y,z)\in r_{\alpha}$ such that $(h(y),h(z))\in\tilde{r}$ then also $(h(x),h(z))\in\tilde{r}$. So, $h$ satisfied all constraints set by $r_{\alpha}]$. Hence, $w\in L_{NE}(\alpha,[r_{\alpha}])$.
($\Leftarrow$): Let $w\in L_{NE}(\alpha,[r_{\alpha}])$ and let $h\in H$ such that $h((\alpha,[r_{\alpha}]))=w$. As $r_{\alpha}\subseteq[r_{\alpha}]$, all constraints in $r_{\alpha}$ must be satisfied. Thus, we also have $w\in L_{NE}(\alpha,r_{\alpha})$. ∎
Following on that, we can also obtain the following relationship.
Lemma 4(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being some equivalence relation and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. If $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, then $\alpha=\beta$.
Using the previous lemma, we can show the following result.
Lemma 5(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns. Then $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$ if and only if $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
Finally, taking all previous lemmas into account, we obtain the main result of this section.
Theorem 5.1(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns. Then $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ if and only if $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
Hence, for the considered relations in this paper, we obtain the following.
Corollary 3(✂)
@@ 5 The Equivalence Problem
Let $R=\{\tilde{r}\}$ for some $\tilde{r}\in\{r_{=},r_{ab},r_{con^{+}}\}$. Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, we can check in polynomial time whether $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
6 Conclusion
We considered the three main decision problems membership, inclusion, and equivalence for relational pattern languages under various typical binary relations, i.e., $r_{=}$, $r_{ssq}$, $r_{ab}$, $_{perm}, rr$$_{rev}$, $r_{com^{+}}$, $r_{com^{*}}$, and $r_{*}$.
In particular, for the membership problem, we observed that in all considered cases, that is, for all relations in $\mathcal{R}$ and for erasing as well as non-erasing relational pattern languages, the problem turns out to be NP-complete, even for very simple relations such as $r_{|w|}$. As the reduction relies on rather simple properties to hold between related areas inside the constructed patterns, we expect the problem to be NP-hard for generally almost all typical relations that could also be considered. This allows for the assumption that searching for words in sets of texts that are compactly described by, e.g., a relational pattern of almost any kind, is generally a computationally hard problem.
For the inclusion problem, we observed that many cases remain undecidable. In the erasing as well as the non-erasing cases, the problem remains undecidable for the structurally less restrictive relations $r_{ssq}$, $r_{ab}$, and $r_{*}$ as well as for the relation $r_{rev}$. In the cases of $r_{ssq}$ and $r_{*}$, we can abuse the property that relating two variables in both directions results basically in $r_{=}$. For the other cases, we observe that the key for the reductions to work is, first, finding ways to force certain variables to be substituted by equal single letters to realize structural properties and, second, to force equality of longer unary words to realize the comparison of counter values. It would be interesting to know whether similar results can be obtained using the other relations that enforce even less structure in valid substitutions.
Regarding the equivalence problem, we obtained decidability in polynomial time for $r_{ab}$ and $r_{com^{+}}$ in the non-erasing case, similar to $r_{=}$. We conjecture that this should hold for many more relations in the non-erasing case by checking some set of minimal length substitutions. For the erasing cases, everything except $r_{|w|}$ remains open and we expect the decidability question to be generally rather difficult to answer, as it is the main open question for $r_{=}$.
This concludes this paper. See Table 3 for a visual summary of the results.
$r_{=}$ $r_{|w|}$ $r_{ssq}$ $r_{ab}$ $_{perm} rr$$_{rev}$ $r_{com^{*}}$ $_{com^{+}} rr_{*}$
 E ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
 E ($\subseteq$)  UD  P UD UD ? UD ? ? UD
 E (==) ?  P ? ? ? ? ? ? ?
 NE ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
 NE ($\subseteq$)  UD  P UD UD ? UD ? ? UD
 NE (==)  P  P ? P ?  D ? P ?
Table 3: Current state of decision problems for relational pattern languages for various constraints (NPC = NP-Complete, UD = Undecidable, D = Decidable, E = Erasing, NE = Non-Erasing, ($\in$) membership problem, ($\subseteq$) inclusion problem, (==) equivalence problem). Red text marks the results covered in this paper.
Appendix 0.A Additional Content for Membership Problem
Proof of Proposition˜1
Proof
Assume $|\Sigma|\geq 2$. Let $r\in\mathcal{R}$, $R=\{r\}$, $w\in\Sigma^{*}$, and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. We can guess a substitution $h\in H$ which results in $|h(\alpha)|=|w|$, then check whether $(x,y)\in r$ for all $(r,x,y)\in r_{\alpha}$, and finally, if that is the case, return true if $h(\alpha)=w$ and return false in any other case. As all relations $r\in\mathcal{R}$ are polynomial-time decidable relations, this results in a polynomial runtime of the verifier, hence resulting for the problem to be in NP. ∎
Proof of Lemma˜1:
Proof
By the result from Angluin in [Angluin1980], we know that this statement already holds for the case of $r_{=}$. This is shown using a reduction from the $3\mathtt{-SAT}$ problem. We quickly recall the construction used in [Angluin1980] and refer to the original source for the proof of its correctness.
Notice that the pattern constructed here is not a relational pattern by our definition but a pattern using the usual notation allowing for multiple occurrences of the same variable. We will comment on the adaptions needed for the notation of relational pattern languages after giving the construction.
Let $\varphi$ be a $3\mathtt{-SAT}$ formula. Let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$. Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$. For each clause $C_{i}\in C$, refer to its three variables by $X_{i,1}$, $X_{i,2}$, and $X_{i,3}$. Define a function $f:\mathcal{X}\rightarrow V$ by $f(X_{i})=v_{i}$ and $f(\overline{X_{i}})=u_{i}$ for new independent pattern variables $u_{i},v_{i}\in V$ with $i\in[m]$. Now, construct the word
$w=\#\ s_{1}\ \#\ ...\ \#\ s_{m}\ \#\ t_{1}\ \#\ ...\ \#\ t_{n}\ \#\ w_{1}\ \#\ ...\ \#\ w_{n}\ \#$
with $s_{i}=1^{3}$, $t_{j}=1^{7}$, and $w_{j}=1^{4}$ for $i\in[m]$ and $j\in[n]$ as well as a pattern
$\alpha=\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\ r_{1}\ \#\ ...\ \#\ r_{n}\ \#$
with $p_{i}=u_{i}v_{i}$, $q_{j}=f(X_{j,1})f(X_{j,2})f(x_{j,3})z_{j}$, and $r_{j}=z_{j}d_{j}$ for new independent pattern variables $z_{j},d_{j}\in X$ with $j\in[n]$. Notice that the construction uses three different elements. The first element contains blocks $u_{i}v_{i}\#$ for $i\in[m]$ which have to be non-emptily substituted to $1^{3}\#$, resulting in one of the variables to be substituted to $11$ and the other one to be substituted to $1$, representing either $X_{i}$ or $\overline{X_{i}}$ to be mapped to true in $\varphi$, having the other one false. The middle part $f(X_{i,1})f(X_{i,2})f(X_{i,3})z_{i}\#$ needs to be non-emptily mapped to $1^{7}$. Notice that due to the last part $z_{i}d_{i}\#$, which is to be substituted to $1^{4}\#$, the variable $z_{i}$ may only be substituted tp $1$, 1111 or $111$, as $d_{i}$ has do be substituted at least to a word of length $1$, i.e., the word $1$. Because of that, at least one of $f(X_{i,j})$ needs to be mapped to $11$ instead of $1$ to obtain $1^{7}\#$, representing at least one variable in each clause to be mapped to true. As this holds for all clauses, deciding whether $w\in L_{NE}(\alpha)$ also decides the satisfiability of the formula $\varphi$. Consider [Angluin1980] for the formal proof.
To obtain this construction in the notation of relational pattern languages, instead of introducing each variable $k\in\mathbb{N}$ times, introduce $k$ different variables in a pattern $\alpha^{\prime}$ that has exactly the same terminal letters and relate those $k$ variables with equality $r_{=}$ in the relation set $r_{\alpha^{\prime}}$. This results in a relational pattern $(\alpha^{\prime},r_{\alpha})$ with $L_{NE}(\alpha)=L_{NE}(\alpha^{\prime},r_{\alpha^{\prime}})$ for which deciding the membership $w\in L_{NE}(\alpha,r_{\alpha})$ also decides the satisfiablity of $\varphi$.
Now we give arguments why the exact same construction works for the other relations in $\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$.
@@ Appendix 0.A Additional Content for Membership Problem
    If $R=\{r_{|w|}\}$, then variables have to be substituted with equal length. As any substitution resulting in $w$ only uses unary words for the variables, having equal length also means being the equal word. This concludes this case.

### v2

@@ 4.1 The Non-Erasing Cases
In the non-erasing case, first notice that, similar to the equivalence problem, in the case of $r_{\varepsilon_{=}}$, we basically compare two regular patterns (as $\varepsilon$ substitutions are excluded anyways). Hence, this case is clearly decidable. Additionally, an interesting observation regarding $r_{ab}$ can be found. Reconstructing the reduction from [5] in the setting of relational pattern languages, it can be shown that the inclusion problem for non-erasing relational pattern languages under $r_{ab}$ or $r_{rev}$ are actually also undecidable, even though $r_{ab}$ generally implies a significantly less restrictive behavior than $r_{=}$ and $r_{rev}$ implies a generally distinct structure.
Theorem 4.1(✂0.B.1).
Let $R:=\{r\}$ for $r\in\{r_{ab},r_{rev}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$. For relational patterns for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$. .
Indeed, this can be shown by reducing the problem of deciding whether a very specific universal Turing machine $U$ has any valid computations to the problem of whether the NE-relational pattern language of one relational pattern under abelian equivalence is included in the NE-relational pattern language of another relational pattern under abelian equivalence. It is done by constructing two such relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ for which we have that $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,\beta_{\alpha})$ if and only if $U$ has no valid computation.
As mentioned above, this result follows from the same reduction as in [5] for the inclusion problem of NE-pattern languages over the equality relation. This is due to the fact that related variables under $r_{ab}$ and $r_{rev}$ result in equal substitutions if only letters or longer but unary words are considered. Hence, the general reasoning of the construction carries over. Only some technical modifications have to be made to formally get it to work in the setting of relational pattern languages under $r_{ab}$ and $r_{rev}$. For the sake of completeness and for verification purposes, an extended sketch of the proof idea can be found in Appendix 0.B.1.
@@ 4.2 The Erasing Cases
We continue with erasing relational pattern languages. We show that the inclusion problem is decidable under $r_{\varepsilon_{=}}$ and undecidable under the reversal relation $r_{rev}$ and under abelian equivalence $r_{ab}$. Indeed, in the case of $r_{\varepsilon_{=}}$, we can use the exact same approach as for the equivalence problem from Theorem 3.2 to obtain decidability in polynomial space using the construction of two NFAs that represent $L_{E}(\alpha,r_{\alpha})$ and $L_{E}(\beta,r_{\beta})$ respectively.
Theorem 4.2.
@@
Let $R=\{r_{\varepsilon_{=}}\}$. Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, checking $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is decidable in polynomial space.
Again, for some pattern structures, we introduce some approaches with improved polynomial running time.
Lemma 4(✂0.B.2).
Let $R=\{r_{\varepsilon_{=}}\}$. Given some alphabet $\Sigma$ and relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
@@
    If $\alpha$ and $\beta$ are terminal-free, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.